# Monotone Lipschitz-Gradient Denoiser: Explainability of Operator Regularization Approaches Free From Lipschitz Constant Control

### v2

@@ Abstract
Index Terms:
weakly convex function, proximity operator, nonexpansive operator, convex optimization
@@ I Introduction
Some promising approaches combining convex and nonconvex functions have been studied [26, 27, 28]. In these approaches, a weakly convex penalty is coupled with a strongly convex fidelity so that convexity of the whole objective function is preserved. A generalized penalty function has been proposed in [29] so that overall convexity can be preserved when applied to inverse problems of underdetermined linear systems (where the fidelity function is not strongly convex). In [30], the penalty has further been generalized to permit a composition with a bounded linear operator, and the generalized penalty has been shown to improve the performance of the $\ell_{1}$-based total variation as a particular instance. In [31], the convex-nonconvex strategy has been applied to outlier-robust signal recovery, and a general framework has been established to clarify when the sufficient condition of overall convexity is also a necessary condition, i.e., when the bound of the regularization parameter is tight. Its extension has also been presented in [32].
@@
In our approach, we suppose that an operator is given. It could be designed by human, or it could also be learned by machine. The designed/learned operator replaces the proximity operator appearing in an operator splitting algorithm, and the optimization problem associated with the resultant algorithm is “derived” subsequently. As such, the processes of “design” and “derivation” are in the reverse order compared to the typical optimization-based approaches. We clarify that the estimate is not explicitly characterized as a minimizer (nor a stationary/critical point) of a user-designed function. It can only be characterized as a minimizer of a convex function involving an implicit regularizer induced by a user-designed “operator”.111 It can also be characterized as a fixed point of some averaged nonexpansive operator from a wider viewpoint. A practical advantage of the proposed approach is that the operator can be designed in a way that it can be computed simply but its corresponding cost function could be complicated, or even impossible to express in a closed form.
@@
where $T$ is a nonlinear operator from a real Hilbert space ${\mathcal{H}}$ to ${\mathcal{H}}$, $f:{\mathcal{H}}\rightarrow{\mathbb{R}}$ is a smooth convex function, and $\mu\in{\mathbb{R}}_{++}$ is the step size. Here, the term smooth (or $\kappa$-smooth more specifically) is used in this paper to mean that the function is differentiable with its gradient Lipschitz continuous with constant $\kappa>0$ (see (3)) over ${\mathcal{H}}$. If in particular $T$ is Moreau’s proximity operator of a proper lower semicontinuous convex function $g$, (1) is the classical forward-backward splitting algorithm to minimize $\mu f+g$. It will turn out that $(x_{k})_{k\in{\mathbb{N}}}$ converges weakly to a minimizer (if exists) of a certain cost function under the following condition (in addition to other technical assumptions).222 One may consider such a condition that $T$ is a selection of the subdifferential of a (not necessarily differentiable) proper lower-semicontinuous convex function $\psi$. However, this seemingly weaker condition together with Lipschitz continuity implies Fréchet differentiability of $\psi$ [39, Proposition 17.41]. To the best of authors’ knowledge, the subdifferentiability is often used in convex analysis to relate an operator to optimization. The case of $\beta\geq 1$ corresponds to Moreau’s proximity operator of convex function [40], thus excluded in the present study.
@@
The term “denoiser” is used in Definition 1 rather than the general term “operator”, because we primarily consider denoising operator (which is often referred to as denoiser in the literature) for $T$. The term “monotone” stems from the fact that convexity of $\psi$ can be characterized by monotonicity of the gradient $\nabla\psi$ (see Section II-D). The convergence argument can be extended to other operator splitting algorithms than (1), as elaborated later on. It should be mentioned that (i) the Lipschitz continuity imposed on MoL-Grad denoiser is weaker than (averaged) nonexpansiveness which requires the Lipschitz constant to be one (or smaller), (ii) every MoL-Grad denoiser is $\beta$-cocoercivity of $T$, i.e., $\beta T$ is firmly nonexpansive, and (iii) the present study includes nonseparable operators in scope.333 For instance, when used as a penalty function for the linear inverse problem with underdetermined systems, a nonconvex function must be nonseparable to preserve overall convexity; i.e., there exists no separable nonconvex function that makes the overall cost function convex [41]. We mention that, when a given denoiser $T:{\mathbb{R}}^{n}\rightarrow{\mathbb{R}}^{n}$ is a continuously differentiable mapping, it can be expressed as $T=\nabla h$ for some $h:{\mathbb{R}}^{n}\rightarrow{\mathbb{R}}$ (i.e., $T$ is a conservative vector field) if and only if $T$ possesses a symmetric Jacobian matrix at every point [42].
@@ I-A Contributions
Given a proper function $f:{\mathcal{H}}\rightarrow(-\infty,+\infty]$, let $T:{\mathcal{H}}\rightarrow{\mathcal{H}}$ be a continuous444 Continuity of the denoiser is often assumed in the literature [43]. operator such that $T(x)\in\operatornamewithlimits{argmin}_{y\in{\mathcal{H}}}[f(y)+(1/2)\left\|x-y\right\|^{2}]\neq\varnothing$ for every $x\in{\mathcal{H}}$. Then, the penalized function $f+(1/2)\left\|x-\cdot\right\|^{2}$ has a unique minimizer for every $x\in{\mathcal{H}}$.
@@
In the second part (Sections IVVI), we consider the modified operator splitting algorithms with the proximity operator replaced by a MoL-Grad denoiser, analyzing the convergence based on the results established in the first part. As a case study, we highlight two specific algorithms: (i) the forward-backward splitting type algorithm given in (1), and (ii) the primal-dual splitting type algorithm which is basically the Condat–Vũ algorithm [8, 9] but with a certain modification. The specific contributions are listed below.
@@ II Preliminaries
[…]
@@ II-B Convexity and Conjugation
Let $f:{\mathcal{H}}\rightarrow(-\infty,+\infty]$ be a proper convex function. Then, we have (i) $\partial f(x)\neq\varnothing$ if $f$ is continuous at $x\in{\mathcal{H}}$ [39, Proposition 16.17], and (ii) $\partial f(x)=\{\nabla f(x)\}$ if $f$ is Gâteaux differentiable, where $\nabla f(x)$ is the Gâteaux gradient of $f$ at $x$555 (Gâteaux and Fréchet gradients of function) Let $U$ be an open subset of ${\mathcal{H}}$. Then, a function $f:U\rightarrow{\mathbb{R}}$ is Gâteaux differentiable at a point $x\in{\mathcal{H}}$ if there exists a vector $\nabla f(x)\in{\mathcal{H}}$ such that $\lim_{\delta\rightarrow 0}\frac{f(x+\delta h)-f(x)}{\delta}=\left\langle{\nabla f(x)},{h}\right\rangle$ for every $h\in{\mathcal{H}}$. On the other hand, a function $f:U\rightarrow{\mathbb{R}}$ is Fréchet differentiable at a point $x\in{\mathcal{H}}$ if there exists a vector $\nabla f(x)\in{\mathcal{H}}$ such that $\lim_{h\rightarrow 0}\frac{f(x+h)-f(x)-\left\langle{\nabla f(x)},{h}\right\rangle}{\left\|h\right\|}=0$. In each case, $\nabla f(x)$ is the Gâteaux (or Fréchet) gradient of $f$ at $x$. If $f$ is Fréchet differentiable over $U$, it is also Gâteaux differentiable over $U$, and the Gâteaux and Fréchet gradients coincide. Those readers who are unfamiliar with convex analysis may consider $\nabla f$ to be the standard gradient without introducing themselves to the general notion of Gâteaux differentiability.
@@ II-C Nonexpansiveness, Cocoercivity, and Fixed Point
    An operator $T:{\mathcal{H}}\rightarrow{\mathcal{H}}$ is continuous (as a mapping from the normed space $({\mathcal{H}},\left\|\cdot\right\|)$ to itself $({\mathcal{H}},\left\|\cdot\right\|)$) if $T$ is continuous at every point $\hat{x}\in{\mathcal{H}}$, i.e., for every sequence $(x_{k})_{k\in{\mathbb{N}}}$, $x_{k}\rightarrow\hat{x}$ implies $T(x_{k})\rightarrow T(\hat{x})$.
@@
Let $T:{\mathcal{H}}\rightarrow{\mathcal{H}}$ be a nonexpansive operator such that ${\rm Fix}\left({T}\right)\neq\varnothing$. Define $T_{\alpha_{k}}:=(1-\alpha_{k}){\rm Id}+\alpha_{k}T$ for a real-number sequence $(\alpha_{k})_{k\in{\mathbb{N}}}\subset[0,1]$ such that $\sum_{k\in{\mathbb{N}}}\alpha_{k}(1-\alpha_{k})=+\infty$.666 It suffices that $\alpha_{k}\in[\epsilon,1-\epsilon]$ ($\forall k\in{\mathbb{N}}$) for some small $\epsilon\in(0,1/2)$. Then, given an initial point $x_{0}\in{\mathcal{H}}$, the sequence $(x_{k})_{k\in{\mathbb{N}}}$ generated by
@@ II-D Monotonicity
$\left\langle{x\!-\!y},{u\!-\!v}\right\rangle\geq 0,~{}\forall(x,u)\in{\rm gra~{}}T,~{}\forall(y,v)\in{\rm gra~{}}T,$ (5)
@@ II-E Proximity Operator and Moreau Envelope
Let $f:{\mathcal{H}}\rightarrow(-\infty,+\infty]$ be a proper function. Given a positive constant $\gamma>0$, assume that $f+(1/2\gamma)\left\|x-\cdot\right\|^{2}$ has a unique minimizer for every fixed $x\in{\mathcal{H}}$. Then, the s-prox operator of $f$ of index $\gamma$ is defined by777 For proper (not necessarily convex) functions, the proximity operator is often defined as a set-valued operator because the $f+(1/2\gamma)\left\|x-\cdot\right\|^{2}$ may have multiple minimizers [45, 46] (see also https://proximity-operator.net/proximityoperator.html). In the present work, we focus on the case in which a unique minimizer exists over the entire space. Although the multi-valued proximity operator coincides with our s-prox operator when a unique minimizer exists, we give the specific name to our operator and we use the different notation ${{\rm s\mbox{-}Prox}}_{\gamma f}$ to convey our message that there are remarkable advantages from the viewpoints of explainability and optimization if we restrict ourselves to the unique-minimizer case.
${{\rm s\mbox{-}Prox}}_{\gamma f}:{\mathcal{H}}\!\rightarrow\!{\mathcal{H}}:x\mapsto\operatornamewithlimits{argmin}_{y\in\mathcal{H}}\!\Big{(}f(y)+\frac{1}{2\gamma}\left\|x\!-\!y\right\|^{2}\Big{)}.$ (6)
@@ III Fundamental Results
The operator ${\rm s\mbox{-}Prox}_{\varphi}$ in (6) is well-defined if $\varphi:=\gamma f$ is $\rho$-weakly convex for $\rho\in(0,1)$ so that $\varphi+(1/2)\left\|x-\cdot\right\|^{2}$ is strongly (and thus strictly) convex and coercive. This simple observation is thoroughly investigated, followed by an extension of Moreau’s decomposition.
@@ III-A Weakly Convex Function and Gradient Operator of Smooth Convex Function
::=:= $~{}\Big{\{}p\in{\mathcal{H}}\mid x\in\partial\Big{(}h+\frac{1}{2}\left\|\cdot\right\|^{2}\Big{)}(p)\Big{\}}.$
@@
$\Leftrightarrow$ $~{}\left\langle{z-y},{s}\right\rangle+J_{x}(y)\leq J_{x}(z),~{}\forall z\in{\mathcal{H}}$
$\Leftrightarrow$ $~{}\left\langle{z-y},{s}\right\rangle+h(y)+\frac{1}{2}\left\|y\right\|^{2}-\left\langle{y},{x}\right\rangle$
@@
$\Leftrightarrow$ $\left\langle{z-y},{s+x}\right\rangle+h(y)+\frac{1}{2}\left\|y\right\|^{2}\leq h(z)+\frac{1}{2}\left\|z\right\|^{2}\!\!\!,~{}\forall z\in{\mathcal{H}}$
$\Leftrightarrow$ $~{}s\in\partial\Big{(}h+\frac{1}{2}\left\|\cdot\right\|^{2}\Big{)}(y)-x,$
@@
$p\in\operatornamewithlimits{argmin}_{y\in{\mathcal{H}}}J_{x}(y)\Leftrightarrow 0\in\partial J_{x}(p)=\partial\Big{(}h+\frac{1}{2}\left\|\cdot\right\|^{2}\Big{)}(p)-x$
$\Leftrightarrow x\in\partial\Big{(}h+\frac{1}{2}\left\|\cdot\right\|^{2}\Big{)}(p)\Leftrightarrow p\in\Big{[}\partial\Big{(}h+\frac{1}{2}\left\|\cdot\right\|^{2}\Big{)}\Big{]}^{-1}(x)$ ,
@@
C1 $\Rightarrow$ C2: For the sake of readability, we present a proof which the reader may follow directly with a minimal use of the known results in convex analysis.888 The proof can be shortened by using [39, Example 22.7] and [39, Example 14.1] to verify (16) and (18). Owing to the convexity of $\check{\varphi}:=\varphi+((1-\beta)/2)\left\|\cdot\right\|^{2}\in\Gamma_{0}({\mathcal{H}})$, we have $T^{-1}={\rm s\mbox{-}Prox}_{\varphi}^{-1}=\partial(\varphi+(1/2)\left\|\cdot\right\|^{2})=\partial(\check{\varphi}+(\beta/2)\left\|\cdot\right\|^{2})=\partial\check{\varphi}+\beta{\rm Id}$ by Lemma 1. Given any fixed $x,y\in{\mathcal{H}}$, let $u\in T^{-1}(x)$ and $v\in T^{-1}(y)$ ($\Leftrightarrow x=T(u),~{}y=T(v)$). Then, since $T^{-1}-\beta{\rm Id}=\partial\check{\varphi}$ is a monotone operator, it follows that
@@
$\Leftrightarrow$ $\left\langle{x-y},{u-v}\right\rangle\geq\beta\left\|x-y\right\|^{2}$
$\Leftrightarrow$ $\left\langle{T(u)-T(v)},{u-v}\right\rangle\geq\beta\left\|T(u)-T(v)\right\|^{2}$
$\Leftrightarrow$ $\left\langle{\beta T(u)-\beta T(v)},{u-v}\right\rangle\geq\left\|\beta T(u)-\beta T(v)\right\|^{2},$ (16)
@@
$[^{\beta}(\check{\varphi}^{*})]^{*}(y)=$ $\sup_{x\in{\mathcal{H}}}\Big{[}\left\langle{x},{y}\right\rangle-\inf_{z\in{\mathcal{H}}}\Big{(}\check{\varphi}^{*}(z)+\frac{1}{2\beta}\left\|z-x\right\|^{2}\Big{)}\Big{]}$
@@
where (${}^{\beta}(\check{\varphi}^{*})$, and thus $\psi$, is a Fréchet differentiable function with $\beta^{-1}$-Lipschitz continuous gradient (see Section II-E):
@@
Theorem 1(b) could be proved using known results as follows. Since $\nabla\psi$ is $\beta^{-1}$-Lipschitz continuous under condition (C2), it follows by [39, Theorem 18.15] that $\psi={}^{\beta}(\check{\varphi}^{*})$, where $\check{\varphi}:=\psi^{*}-(\beta/2)\left\|\cdot\right\|^{2}\in\Gamma_{0}({\mathcal{H}})$. Hence, by Moreau’s decomposition, it follows that $\nabla\psi=\beta^{-1}({\rm Id}-{\rm s\mbox{-}Prox}_{\beta\check{\varphi}^{*}})={\rm s\mbox{-}Prox}_{\beta^{-1}\check{\varphi}}\circ(\beta^{-1}{\rm Id})={\rm s\mbox{-}Prox}_{\beta^{-1}\psi^{*}-(1/2)\left\|\cdot\right\|^{2}}\circ(\beta^{-1}{\rm Id})={\rm s\mbox{-}Prox}_{\psi^{*}-(1/2)\left\|\cdot\right\|^{2}},$ where the last inequality can be verified by the definition of proximity operator with some manipulations. Here, it is clear that $\varphi:=\psi^{*}-(1/2)\left\|\cdot\right\|^{2}=\check{\varphi}-((1-\beta)/2)\left\|\cdot\right\|^{2}$ is $(1-\beta)$-weakly convex.

### v3

@@ I Introduction
Some promising approaches combining convex and nonconvex functions have been studied [26, 27, 28]. In these approaches, a weakly convex penalty is coupled with a strongly convex fidelity so that the overall convexity of the whole objective function is preserved. A generalized penalty function has been proposed in [29] so that the overall convexity can be preserved even when applied to inverse problems of underdetermined linear systems (where the fidelity function is not strongly convex). In [30], the penalty has further been generalized to permit a composition with a bounded linear operator, and the generalized penalty has been shown to improve the performance of the $\ell_{1}$-based total variation as a particular instance. In [31], the convex-nonconvex strategy has been applied to outlier-robust signal recovery, and a general framework has been established to clarify when the sufficient condition of overall convexity is also a necessary condition, i.e., when the bound of the regularization parameter is tight. Its extension has also been presented in [32].
@@
In our approach, we suppose that an operator is given. It could be designed by human, or it could also be learned by machine. The designed/learned operator replaces the proximity operator appearing in an operator splitting algorithm, and the optimization problem associated with the resultant algorithm is “derived” subsequently. As such, the processes of “design” and “derivation” are in the reverse order compared to the typical optimization-based approaches. We clarify that the estimate is not explicitly characterized as a minimizer (nor a stationary/critical point) of a user-designed function. It can only be characterized as a minimizer of a convex function involving an implicit regularizer induced by a user-designed “operator”. [^1] It can also be characterized as a fixed point of some averaged nonexpansive operator from a wider viewpoint. A practical advantage of the proposed approach is that the operator can be designed in a way that it can be computed simply but its corresponding cost function could be complicated, or even impossible to express in a closed form.
@@
where $T$ is a nonlinear operator from a real Hilbert space $\mathcal{H}$} to $\mathcal{H}$}, $f:{\mathcal{H}}\rightarrow{\mathbb{R}}$ is a smooth convex function, and $\mu\in{\mathbb{R}}_{++}$ is the step size. Here, the term smooth (or $\kappa$-smooth more specifically) is used in this paper to mean that the function is differentiable with its gradient Lipschitz continuous with constant $\kappa>0$ (see (3)) over $\mathcal{H}$}. If in particular $T$ is Moreau’s proximity operator of a proper lower semicontinuous convex function $g$, (1) is the classical forward-backward splitting algorithm to minimize $\mu f+g$. It will turn out that $(x_{k})_{k\in{\mathbb{N}}}$ converges weakly to a minimizer (if exists) of a certain cost function under the following condition (in addition to other technical assumptions). [^2] One may consider such a condition that $T$ is a selection of the subdifferential of a (not necessarily differentiable) proper lower-semicontinuous convex function $\psi$. However, this seemingly weaker condition together with (Lipschitz) continuity implies Fréchet differentiability of $\psi$ [39, Proposition 17.41]. To the best of authors’ knowledge, the subdifferentiability is often used in convex analysis to relate an operator to optimization. The case of $\beta\geq 1$ corresponds to Moreau’s proximity operator of convex function [40], thus excluded in the present study.
@@
The term “denoiser” is used in Definition 1 rather than the general term “operator”, because we primarily consider denoising operator (which is often referred to as denoiser in the literature) for $T$. The term “monotone” stems from the fact that convexity of $\psi$ can be characterized by monotonicity of the gradient $\nabla\psi$ (see Section II-D). The convergence argument can be extended to other operator splitting algorithms than (1), as elaborated later on. It should be mentioned that (i) the Lipschitz continuity imposed on MoL-Grad denoiser is weaker than (averaged) nonexpansiveness which requires the Lipschitz constant to be one (or smaller), (ii) every MoL-Grad denoiser is $\beta$-cocoercivity of $T$, i.e., $\beta T$ is firmly nonexpansive, and (iii) the present study includes nonseparable operators in scope. [^3] For instance, when used as a penalty function for the linear inverse problem with underdetermined systems, a nonconvex function must be nonseparable to preserve overall convexity; i.e., there exists no separable nonconvex function that makes the overall cost function convex [41]. We mention that, when a given denoiser $T:{\mathbb{R}}^{n}\rightarrow{\mathbb{R}}^{n}$ is a continuously differentiable mapping, it can be expressed as $T=\nabla h$ for some $h:{\mathbb{R}}^{n}\rightarrow{\mathbb{R}}$ (i.e., $T$ is a conservative vector field) if and only if $T$ possesses a symmetric Jacobian matrix at every point [42].
@@ I-A Contributions
Given a proper function $f:{\mathcal{H}}\rightarrow(-\infty,+\infty]$, let $T:{\mathcal{H}}\rightarrow{\mathcal{H}}$ be a continuous [^4] Continuity of the denoiser is often assumed in the literature [43]. operator such that $T(x)\in\operatornamewithlimits{argmin}_{y\in{\mathcal{H}}}[f(y)+(1/2)\left\|x-y\right\|^{2}]\neq\varnothing$ for every $x\in{\mathcal{H}}$. Then, the penalized function $f+(1/2)\left\|x-\cdot\right\|^{2}$ has a unique minimizer for every $x\in{\mathcal{H}}$.
@@
In the second part (Sections IVVI), we consider the modified operator splitting algorithms with the proximity operator replaced by a MoL-Grad denoiser, analyzing the convergence based on the results established in the first part. As a case study, we highlight two specific algorithms: (i) the forward-backward splitting type algorithm given in (1), and (ii) the primal-dual splitting type algorithm which is basically the Condat–Vũ algorithm [8, 9]. The specific contributions are listed below.
@@ II Preliminaries
Let $({\mathcal{H}},\left\langle{\cdot},{\cdot}\right\rangle)$ be a real Hilbert space with the induced norm $\left\|\cdot\right\|$. Let $\mathbb{R}$}, $\mathbb{R}$$_{+}$, $\mathbb{R}$$_{++}$, and $\mathbb{N}$} denote the sets of real numbers, nonnegative real numbers, strictly positive real numbers, and nonnegative integers, respectively.
@@ II-B Convexity and Conjugation
Let $f:{\mathcal{H}}\rightarrow(-\infty,+\infty]$ be a proper convex function. Then, we have (i) $\partial f(x)\neq\varnothing$ if $f$ is continuous at $x\in{\mathcal{H}}$ [39, Proposition 16.17], and (ii) $\partial f(x)=\{\nabla f(x)\}$ if $f$ is Gâteaux differentiable, where $\nabla f(x)$ is the Gâteaux gradient of $f$ at xx.555(Gâteaux and Fréchet gradients of function) Let $U$ be an open subset of $\mathcal{H}$}. Then, a function $f:U\rightarrow{\mathbb{R}}$ is Gâteaux differentiable at a point $x\in{\mathcal{H}}$ if there exists a vector $\nabla f(x)\in{\mathcal{H}}$ such that $\lim_{\delta\rightarrow 0}\frac{f(x+\delta h)-f(x)}{\delta}=\left\langle{\nabla f(x)},{h}\right\rangle$ for every $h\in{\mathcal{H}}$. On the other hand, a function $f:U\rightarrow{\mathbb{R}}$ is Fréchet differentiable at a point $x\in{\mathcal{H}}$ if there exists a vector $\nabla f(x)\in{\mathcal{H}}$ such that $\lim_{h\rightarrow 0}\frac{f(x+h)-f(x)-\left\langle{\nabla f(x)},{h}\right\rangle}{\left\|h\right\|}=0$. In each case, $\nabla f(x)$ is the Gâteaux (or Fréchet) gradient of $f$ at $x$. If $f$ is Fréchet differentiable over $U$, it is also Gâteaux differentiable over $U$, and the Gâteaux and Fréchet gradients coincide. Those readers who are unfamiliar with convex analysis may consider $\nabla f$ to be the standard gradient without introducing themselves to the general notion of Gâteaux differentiability.
@@ II-C Nonexpansiveness, Cocoercivity, and Fixed Point
[…]
@@
Let $T:{\mathcal{H}}\rightarrow{\mathcal{H}}$ be a nonexpansive operator such that ${\rm Fix}\left({T}\right)\neq\varnothing$. Define $T_{\alpha_{k}}:=(1-\alpha_{k}){\rm Id}+\alpha_{k}T$ for a real-number sequence $(\alpha_{k})_{k\in{\mathbb{N}}}\subset[0,1]$ such that $\sum_{k\in{\mathbb{N}}}\alpha_{k}(1-\alpha_{k})=+\infty$. [^6] It suffices that $\alpha_{k}\in[\epsilon,1-\epsilon]$ ($\forall k\in{\mathbb{N}}$) for some small $\epsilon\in(0,1/2)$. Then, given an initial point $x_{0}\in{\mathcal{H}}$, the sequence $(x_{k})_{k\in{\mathbb{N}}}$ generated by
@@ II-D Monotonicity
$\hskip-20.00003pt\left\langle{x\!-\!y},{u\!-\!v}\right\rangle\geq 0,~\forall(x,u)\in{\rm gra~}T,~\forall(y,v)\in{\rm gra~}T,$ (5)
@@ II-E Proximity Operator and Moreau Envelope
Let $f:{\mathcal{H}}\rightarrow(-\infty,+\infty]$ be a proper function. Given a positive constant $\gamma>0$, assume that $f+(1/(2\gamma))\left\|x-\cdot\right\|^{2}$ has a unique minimizer for every fixed $x\in{\mathcal{H}}$. Then, the s-prox operator of $f$ of index $\gamma$ is defined by [^7] For proper (not necessarily convex) functions, the proximity operator is often defined as a set-valued operator because the $f+(1/(2\gamma))\left\|x-\cdot\right\|^{2}$ may have multiple minimizers [45, 46] (see also https://proximity-operator.net/proximityoperator.html). In the present work, we focus on the case in which a unique minimizer exists over the entire space. Although the multi-valued proximity operator coincides with our s-prox operator when a unique minimizer exists, we give the specific name to our operator and we use the different notation ${{\rm s\mbox{-}Prox}}_{\gamma f}$ to convey our message that there are remarkable advantages from the viewpoints of explainability and optimization if we restrict ourselves to the unique-minimizer case.
$\hskip-13.00005pt{{\rm s\mbox{-}Prox}}_{\gamma f}:{\mathcal{H}}\!\rightarrow\!{\mathcal{H}}:x\mapsto\operatornamewithlimits{argmin}_{y\in\mathcal{H}}\!\Big(f(y)+\frac{1}{2\gamma}\left\|x\!-\!y\right\|^{2}\Big).$ (6)
@@ III Fundamental Results
The operator ${\rm s\mbox{-}Prox}_{\gamma f}$ in (6) is well-defined if $\gamma f$ is $\rho$-weakly convex for $\rho\in(0,1)$ so that $\gamma f+(1/2)\left\|x-\cdot\right\|^{2}$ is strongly (and thus strictly) convex and coercive. This simple observation is thoroughly investigated, followed by an extension of Moreau’s decomposition.
@@ III-A Weakly Convex Function and Gradient Operator of Smooth Convex Function
$:=$ $~\Big\{p\in{\mathcal{H}}\mid x\in\partial\Big(h+\frac{1}{2}\left\|\cdot\right\|^{2}\Big)(p)\Big\}.$
@@
$\hskip-20.00003pt\Leftrightarrow$ $~\left\langle{z-y},{s}\right\rangle+J_{x}(y)\leq J_{x}(z),~\forall z\in{\mathcal{H}}$
$\hskip-20.00003pt\Leftrightarrow$ $~\left\langle{z-y},{s}\right\rangle+h(y)+\frac{1}{2}\left\|y\right\|^{2}-\left\langle{y},{x}\right\rangle$
@@
$\hskip-20.00003pt\Leftrightarrow$ $\left\langle{z-y},{s+x}\right\rangle+h(y)+\frac{1}{2}\left\|y\right\|^{2}\leq h(z)+\frac{1}{2}\left\|z\right\|^{2}\!\!\!,~\forall z\in{\mathcal{H}}$
$\hskip-20.00003pt\Leftrightarrow$ $~s\in\partial\Big(h+\frac{1}{2}\left\|\cdot\right\|^{2}\Big)(y)-x,$
@@
$\hskip-16.00008ptp\in\operatornamewithlimits{argmin}_{y\in{\mathcal{H}}}J_{x}(y)\Leftrightarrow 0\in\partial J_{x}(p)=\partial\Big(h+\frac{1}{2}\left\|\cdot\right\|^{2}\Big)(p)-x$
$\hskip-16.99998pt\Leftrightarrow x\in\partial\Big(h+\frac{1}{2}\left\|\cdot\right\|^{2}\Big)(p)\Leftrightarrow p\in\Big[\partial\Big(h+\frac{1}{2}\left\|\cdot\right\|^{2}\Big)\Big]^{-1}(x)$ ,
@@
C1 $\Rightarrow$ C2: For the sake of readability, we present a proof which the reader may follow directly with a minimal use of the known results in convex analysis. [^8] The proof can be shortened by using [39, Example 22.7] and [39, Example 14.1] to verify (16) and (18). Owing to the convexity of $\check{\varphi}:=\varphi+((1-\beta)/2)\left\|\cdot\right\|^{2}\in\Gamma_{0}({\mathcal{H}})$, we have $T^{-1}={\rm s\mbox{-}Prox}_{\varphi}^{-1}=\partial(\varphi+(1/2)\left\|\cdot\right\|^{2})=\partial(\check{\varphi}+(\beta/2)\left\|\cdot\right\|^{2})=\partial\check{\varphi}+\beta{\rm Id}$ by Lemma 1. Given any fixed $x,y\in{\mathcal{H}}$, let $u\in T^{-1}(x)$ and $v\in T^{-1}(y)$ ($\Leftrightarrow x=T(u),~y=T(v)$). Then, since $T^{-1}-\beta{\rm Id}=\partial\check{\varphi}$ is a monotone operator, it follows that
@@
$\hskip-10.00002pt\Leftrightarrow$ $\left\langle{x-y},{u-v}\right\rangle\geq\beta\left\|x-y\right\|^{2}$
$\hskip-10.00002pt\Leftrightarrow$ $\left\langle{T(u)-T(v)},{u-v}\right\rangle\geq\beta\left\|T(u)-T(v)\right\|^{2}$
$\hskip-10.00002pt\Leftrightarrow$ $\left\langle{\beta T(u)-\beta T(v)},{u-v}\right\rangle\geq\left\|\beta T(u)-\beta T(v)\right\|^{2},$ (16)
@@
$\hskip-20.00003pt[^{\beta}(\check{\varphi}^{*})]^{*}(y)=$ $\sup_{x\in{\mathcal{H}}}\Big[\left\langle{x},{y}\right\rangle-\inf_{z\in{\mathcal{H}}}\Big(\check{\varphi}^{*}(z)+\frac{1}{2\beta}\left\|z-x\right\|^{2}\Big)\Big]$
@@
[…]
@@
[…]